OpenStudy (anonymous):

The monthly cost of driving a car depends on the number of miles driven. Lynn found that in May her driving cost was $380 for 300 mi and in June her cost was $500 for 900 mi. Assume that there is a linear relationship between the monthly cost C of driving a car and the distance driven d. (a) Find a linear equation that relates C and d. (b) Use part (a) to predict the cost of driving 1600 mi per month.

OpenStudy (anonymous):

You have two points May(380,300) and June (500,900). \(\text{Slope} = \dfrac{\text{rise}}{\text{run}} = \dfrac{\text{change in y}}{\text{change in x}} = \dfrac{y_2-y_1}{x_2-x_1}=m\) Point-Slope Formula: \((y-y_1) = m(x-x_1)\)

OpenStudy (anonymous):

so how do i do that

OpenStudy (anonymous):

i got (y-300)=5(x-380)